# Supertranslation ambiguity in post-Minkowskian expansion

### v1

@@
Abstract: The supertranslation ambiguity of angular momentum is a long-standing problem in general relativity. In the context of post-Minkowskian expansion, the supertranslation ambiguity arises the puzzle of the angular momentum loss to start at ${\cal O}(G^{2})$ or at ${\cal O}(G^{3})$ in gravitational scattering. In this paper, we propose a generic prescription to fix the supertranslation ambiguity at the linear order in post-Minkowskian expansion which will uniquely determine the angular momentum loss.
@@ 1 Introduction
A surprising result from Bondi’s framework [3, 4] is that the asymptotic symmetry group at null infinity is not the Poincaré group, but the semidirect sum of supertranslations, an infinite dimensional Abelian group which extends the translations, with the global conformal group of the celestial sphere that is isomorphic to the Lorentz group. Though, the enhancement from translation to supertranslation has several remarkable applications recently, see, e.g., the comprehensive review in [5], it also brings puzzles in general relativity. A long-standing problem is the supertranslation ambiguity in defining angular momentum [6, 7], see also some recent progresses on this issue [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. In the context of the PM expansion, the supertranslation ambiguity causes the puzzle of angular momentum loss in gravitational scattering [21]. The leading order of the angular momentum loss can be at ${\cal O}(G^{2})$ or at ${\cal O}(G^{3})$ depending on different gauge choices under the supertranslation [22, 23, 24, 25, 26, 27, 28, 21, 29, 30, 31, 32, 33, 34]. A self-contained criteria to fix the gauge choice, to the best of our knowledge, is still an open question and a definite angular momentum loss formula is certainly of utmost importance for higher precision PM computations.
@@
In this paper, we propose a generic prescription to fix the supertranslation ambiguity at the linear order in the PM expansion which can uniquely determine the angular momentum loss. The resolution is encoded in the linear nature of the PM expansion. In full Einstein theory, the supertranslation freedom is intrinsic in the sense that it is associated with the propagating degree of freedom. One can not fix the gauge choice associated to the supertranslation freedom. Because gravitational radiation will change the gauge choice in any case [7], see also for a modern interpretation [35]. When the Einstein equation is expanded about the Newton’s constant $G$, one can recover a linearized theory at each order with its own gauge structure which is inherited from the diffeomorphism invariance of the Einstein theory. When one also expands a generic transformation of coordinates system in $G$, the transformation at each order of $G$ is the gauge transformation of the linearized theory at the corresponding order. Adapting the Bondi’s framework to the PM expansion, one can recover a supertranslation at each order of $G$. A fine tuning structure arises during the linearization. The propagating degree of freedom is at ${\cal O}(G^{2})$. While the disagreement of angular momentum loss is relevant to the first order supertranslation. This is a conceptually crucial feature since we can fix the first order supertranslation to resolve the angular momentum loss puzzle without touching the propagating degree of freedom at ${\cal O}(G^{2})$.
@@
Consider that the spacetime is sourced by a system of pointlike bodies. The relativistic definitions of 4-momentum and angular momentum from the stress tensor are the same as the ADM definitions of the linear order metric [21]. Though, the linear metric in the Bondi gauge has first order supertranslation dependence [21, 36], the 4-momentum and angular momentum defined at null infinity are independent of the first order supertranslation and coincide with the relativistic definitions [36]. It is important to point out that the zeroth order supertranslation was not involved at all in [21, 36]. One can think of that the coincidence of the charges at null infinity and the relativistic definition of charges in the bulk rules out the zeroth order supertranslation. Then the key idea we will implement is to use the charges at the second order to control the first order supertranslation. Naively, this is impossible since there is no conservation law at the second order due to the gravitational radiation. Nevertheless, we can consider part of the second order solution. The source of the linearized equations of motion at the second order consists of two parts. The corrected stress tensor of the pointlike bodies due to the linear order metric and the effective stress tensor from the first order metric. Since the theory is linear, we can study those two parts independently. If one ignores the modification of the motion of the pointlike bodies and considers only the effective stress tensor, which we would refer to as the self-part, the relativistic definition of 4-momentum and angular momentum will not be corrected from the next-to-leading order in the expansion of $G$. Hence, it is reasonable to require that the second order metric of the self-part at null infinity should have vanishing 4-momentum and angular momentum. This is the criteria we propose to fix the supertranslation ambiguity at the first order.
@@
Since the waveform at the second order from the source of arbitrary numbers of pointlike bodies is not yet known, nor the self-part of this system, we can not test our prescription in the generic case. We will use an alternative approach to prove that the first order supertranslation freedom will be ruled out from our prescription. We apply the asymptotic analysis to derive the generic second order solution with the effective stress tensor in series expansion near null infinity. Then the criteria of vanishing of 4-momentum and angular momentum will fix the initial data of the second order asymptotic solution. We then compute the transformation laws of the 4-momentum and angular momentum under the first order supertranslation. We show that a generic first order supertranslation will change the next-to-leading order angular momentum. The vanishing conditions of 4-momentum and angular momentum finally fix the first order supertranslation.
@@ 2 Self-part at the second order
where the indices are now raised by $\eta^{\mu\nu}$. The equations of motion for the self-part at the second order are [37] (see also Appendix A)
@@
$T^{E}_{\mu\nu}=\eta_{\mu\nu}\bigg{[}H^{\alpha\beta}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}H+\bar{\nabla}^{2}H_{\alpha\beta}-2\bar{\nabla}^{\tau}\bar{\nabla}_{\alpha}H_{\tau\beta}\right)-\frac{1}{2}\bar{\nabla}_{\beta}H_{\alpha\tau}\bar{\nabla}^{\tau}H^{\alpha\beta}\\ +\bar{\nabla}_{\beta}H\left(\bar{\nabla}_{\alpha}H^{\alpha\beta}-\frac{1}{4}\bar{\nabla}^{\beta}H\right)-\bar{\nabla}_{\alpha}H^{\alpha\beta}\bar{\nabla}^{\tau}H_{\tau\beta}+\frac{3}{4}\bar{\nabla}_{\alpha}H^{\tau\beta}\bar{\nabla}^{\alpha}H_{\tau^{\beta}}\bigg{]}\\ H^{\tau\alpha}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\mu}H_{\nu\tau}+\bar{\nabla}_{\alpha}\bar{\nabla}_{\nu}H_{\mu\tau}-\bar{\nabla}_{\alpha}\bar{\nabla}_{\tau}H_{\mu\nu}-\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}H_{\tau\alpha}\right)\\ +H_{\mu\nu}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}H^{\alpha\beta}-\bar{\nabla}^{2}H\right)+\bar{\nabla}_{\alpha}H^{\alpha\tau}\left(\bar{\nabla}_{\mu}H_{\nu\tau}+\bar{\nabla}_{\nu}H_{\mu\tau}-\bar{\nabla}_{\tau}H_{\mu\nu}\right)\\ -\frac{1}{2}\bar{\nabla}_{\mu}H_{\tau\alpha}\bar{\nabla}_{\nu}H^{\tau\alpha}+\bar{\nabla}_{\alpha}H_{\nu\tau}\bar{\nabla}^{\tau}H_{\mu}^{\alpha}-\bar{\nabla}_{\alpha}H_{\nu\tau}\bar{\nabla}^{\alpha}H_{\mu}^{\tau}\\ -\frac{1}{2}\bar{\nabla}^{\alpha}H\left(\bar{\nabla}_{\mu}H_{\nu\alpha}+\bar{\nabla}_{\nu}H_{\mu\alpha}-\bar{\nabla}_{\alpha}H_{\mu\nu}\right).$ (3)
@@ 3 Solution space at the second order
$\partial_{u}n_{A}=\frac{1}{4}D^{B}D_{A}D^{E}c_{EB}-\frac{1}{4}D^{2}D^{B}c_{BA}-uD_{A}(\partial_{u}m_{b})\\ \frac{1}{4}D_{A}(C_{BE}N^{BE})-\frac{1}{4}D_{B}(C^{BE}N_{EA})+\frac{1}{2}C_{AB}D_{E}N^{EB},$ (18)
@@ 5 First order supertranslations
${r^{\prime}}=r-\frac{G}{2}D^{2}T+\frac{G^{2}}{4r}\bigg{(}2D_{A}C^{AB}D_{B}T+C^{AB}D_{A}D_{B}T-D_{A}D^{2}TD^{A}T$
$\quad\quad\quad-D_{A}D_{B}TD^{A}D^{B}T+\partial_{A}D^{B}T\partial_{B}D^{A}T\bigg{)}+\frac{G^{2}}{4r}\partial_{A}\bigg{(}\gamma^{AB}\partial_{B}D^{C}TD_{C}T$
$\quad\quad\quad-\partial_{C}\gamma^{AB}D_{B}TD^{C}T\bigg{)}+\frac{G^{2}}{4r}\bigg{(}\gamma^{AB}\partial_{B}D^{C}TD_{C}T-\partial_{C}\gamma^{AB}D_{B}TD^{C}T\bigg{)}\partial_{A}\ln\gamma$
$\quad\quad\quad-\frac{G^{2}}{4r}D^{A}TD^{B}T\partial_{A}\partial_{B}\ln{\gamma}+\frac{G^{2}}{6r^{2}}D^{AB}D_{A}D_{B}T+{\cal O}(G^{3}),$ (33)
@@
when deriving (37). Clearly, the condition of preserving $n_{A}$ in the form of (26) in the new coordinates system will reduce the supertranslation to one translation along time direction $u$, considering that $M_{b}=-\sum{\frac{m}{(v_{\mu}n^{\mu})^{3}}}$ can not be a constant for arbitrary numbers of pointlike bodies.
@@ 6 Concluding remarks
In this paper, we apply the charge conservation to prove that the first order supertranslation should be ruled out in the PM expansion which can serve as the foundation of uniquely determining the angular momentum loss in gravitational scattering. It is worthwhile to point out that ruling out the first order supertranslation only means that the first order supertranslation parameter is fixed, but not necessary to be zero. The precise formula of the first order supertranslation parameter can be obtained for a precise gravitational scattering, such as gravitational Bremsstrahlung [48, 23]. It is amusing to notice that once the first order supertranslation is fixed, the complete first order metric is fixed. The derivation of the other parts of the second order metric are relevant to the first order metric, but can not affect the first order metric. Then the key point of our prescription is that we use part of the higher order solutions in $G$ expansion to fix the lower order supertranslation. It is interesting to see if such idea can be valid at higher order in $G$ to fix the complete supertranslation freedom, e.g., for the gravitational Bremsstrahlung.
@@
The first piece on the right hand side is the independent second order gauge transformation. The second piece is induced from the first order gauge transformation. It is not a pure gauge transformation at the second order because it will change the first order metric. One can effectively consider that the second order perturbative theory is living on a curved background with metric $\eta_{\mu\nu}+GH_{\mu\nu}$. The effective stress tensor coupled to the second order theory is from curved effect of the background spacetime. Then the first order gauge transformation does change the spacetime metric, hence change the frame of the observer. Note that the translations along spatial dimensions correspond to the shift of the origin of the spacetime which has also been ruled out as they will change the angular momentum. Thus, a first order supertranslation (except the translation along $u$) will set the the second order theory to a different frame or a different origin choice than the first order theory. Such interpretation is in consistent with the full Einstein theory [16] where memory effect will change the frame of observer which leads to the supertranslation ambiguity of angular momentum. In the PM expansion, there is no dissipation from the self-part at the second order. The charge conservation can serve as calibration of frames and origin choices of the two observers for describing the first order and second order perturbative theories. The fact that theories at different orders can be in different frames or with different origin choices reflects the non-linear effect from interactions between the gravitational fields and the observers in the full Einstein theory.
@@ Acknowledgments
The authors thank Zhengwen Liu and Kai-Yu Zhang for useful discussions. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grants No. 11935009 and No. 11905156.
@@ Appendix B Conservation of the effective stress tensor
The conservation of the stress tensor is induced from the Bianchi identity of the Einstein tensor $\nabla_{\mu}G^{\mu}_{\nu}=0$. We consider the Minkowski background and the pointlike bodies source for the first order perturbative theory. Hence the zeroth order of the Einstein tensor vanishes and the first order vanishes outside the trajectory of the pointlike bodies. At the second order, the Bianchi identity is reduced to
@@
It is obvious that ¯$\bar{\nabla}^{\mu}E_{\mu\nu}=0$ since the linearized equation $E_{\mu\nu}$ at the second order is invariant under the gauge transformation

### v2

@@
Abstract: The supertranslation ambiguity of angular momentum is a long-standing problem in general relativity. In the context of post-Minkowskian expansion, the supertranslation ambiguity arises the puzzle of the angular momentum loss to start at ${\cal O}(G^{2})$ or at ${\cal O}(G^{3})$ in gravitational scattering. In this paper, we propose a generic prescription to fix the supertranslation ambiguity at the linear order in post-Minkowskian expansion which will uniquely determine the angular momentum loss. The proposal is self-contained in the post-Minkowskian expansion and involves only null infinity data.
@@ 1 Introduction
A surprising result from Bondi’s framework [3, 4] is that the asymptotic symmetry group at null infinity is not the Poincaré group, but the semidirect sum of supertranslations, an infinite dimensional Abelian group which extends the translations, with the global conformal group of the celestial sphere that is isomorphic to the Lorentz group. Though, the enhancement from translation to supertranslation has several remarkable applications recently, see, e.g., the comprehensive review in [5], it also brings puzzles in general relativity. A long-standing problem is the supertranslation ambiguity in defining angular momentum [6, 7]. In the context of the PM expansion, the supertranslation ambiguity causes the puzzle of angular momentum loss in gravitational scattering [8]. The leading order of the angular momentum loss can be at ${\cal O}(G^{2})$ or at ${\cal O}(G^{3})$ depending on different gauge choices under the supertranslation [9, 10, 11, 12, 13, 14, 15, 8, 16, 17, 18, 19, 20, 21, 22]. Angular momentum loss formula at null infinity and its relation to the change of angular momentum of the source are certainly of utmost importance for higher precision PM computations, which lead to the recent investigations from various perspectives, including the binary systems [23, 24], supertranslation invariant definition of angular momentum at null infinity [25, 26, 27, 28, 29, 30, 31] and at spatial infinity [32, 33], supertranslation frame-fixing at the far past of the future null infinity [8, 34], supertranslation frame-fixing from covariance [35], memory effect [36, 37], and flux balance law [37, 38].
@@
In this paper, we propose a self-contained prescription in the PM expansion to fix the supertranslation ambiguity at the linear order which can uniquely determine the angular momentum loss. The resolution is encoded in the linear nature of the PM expansion and involves only null infinity data. In Einstein theory, the supertranslation freedom is intrinsic in the sense that it is associated with the propagating degree of freedom. One can not fix the gauge choice associated to the supertranslation freedom. Because gravitational radiation will change the gauge choice in any case [7], see also [39] for a modern interpretation. When the Einstein equation is expanded about the Newton’s constant $G$, one can recover a linearized theory at each order with its own gauge structure which is inherited from the diffeomorphism invariance of the Einstein theory. When one also expands a generic transformation of coordinates system in $G$, the transformation at each order of $G$ is the gauge transformation of the linearized theory at the corresponding order. Adapting the Bondi’s framework to the PM expansion, one can recover a supertranslation at each order of $G$. A fine tuning structure arises during the linearization. The propagating degree of freedom is at ${\cal O}(G^{2})$. While the disagreement of angular momentum loss is relevant to the first order supertranslation. This is a conceptually crucial feature since we can fix the first order supertranslation to resolve the angular momentum loss puzzle without touching the propagating degree of freedom at ${\cal O}(G^{2})$.
@@
Consider that the spacetime is sourced by a system of pointlike bodies. The relativistic definitions of 4-momentum and angular momentum from the stress tensor are the same as the ADM definitions of the linear order metric [8]. Though, the linear metric in the Bondi gauge has first order supertranslation dependence [8, 40], the 4-momentum and angular momentum defined at null infinity are independent of the first order supertranslation and coincide with the relativistic definitions [40]. It is important to point out that the zeroth order supertranslation was not involved at all in [8, 40]. One can think of that the coincidence of the charges at null infinity and the relativistic definition of charges in the bulk rules out the zeroth order supertranslation. Then the key idea we will implement is to use the charges at the second order to control the first order supertranslation. Naively, this is impossible since there is no conservation law at the second order due to the gravitational radiation. Nevertheless, we can consider part of the second order solution. The source of the linearized equations of motion at the second order consists of two parts. The corrected stress tensor of the pointlike bodies due to the linear order metric and the effective stress tensor from the first order metric. Since the theory is linear, we can study those two parts independently. If one ignores the modification of the motion of the pointlike bodies and considers only the effective stress tensor, which we would refer to as the self-part, [^1] The self-part in our terminology includes the self-terms and cross-terms in a two pointlike objects system investigated in [41]. the relativistic definition of 4-momentum and angular momentum will not be corrected from the next-to-leading order in the expansion of $G$. Hence, it is reasonable to require that the second order metric of the self-part at null infinity should have vanishing 4-momentum and angular momentum. This is the criterion we propose to fix the supertranslation ambiguity at the first order.
@@
Since the waveform at the second order from the source of arbitrary numbers of pointlike bodies is not yet known, nor the self-part of this system, we can not test our prescription in the generic case. We will use an alternative approach to prove that the first order supertranslation freedom will be ruled out by our prescription. We apply the asymptotic analysis to derive the generic second order solution with the effective stress tensor in series expansion near null infinity. Then the criteria of vanishing of 4-momentum and angular momentum will fix the initial data of the second order asymptotic solution. We then compute the transformation laws of the 4-momentum and angular momentum under the first order supertranslation. We show that a generic first order supertranslation will change the next-to-leading order angular momentum. The vanishing conditions of 4-momentum and angular momentum finally fix the first order supertranslation.
@@ 2 Self-part at the second order
where the indices are now raised by $\eta^{\mu\nu}$. The equations of motion for the self-part at the second order are [42, 41] (see also Appendix A)
@@
$T^{E}_{\mu\nu}=\eta_{\mu\nu}\bigg{[}H^{\alpha\beta}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}H+\bar{\nabla}^{2}H_{\alpha\beta}-2\bar{\nabla}^{\tau}\bar{\nabla}_{\alpha}H_{\tau\beta}\right)-\frac{1}{2}\bar{\nabla}_{\beta}H_{\alpha\tau}\bar{\nabla}^{\tau}H^{\alpha\beta}\\ +\bar{\nabla}_{\beta}H\left(\bar{\nabla}_{\alpha}H^{\alpha\beta}-\frac{1}{4}\bar{\nabla}^{\beta}H\right)-\bar{\nabla}_{\alpha}H^{\alpha\beta}\bar{\nabla}^{\tau}H_{\tau\beta}+\frac{3}{4}\bar{\nabla}_{\alpha}H^{\tau\beta}\bar{\nabla}^{\alpha}H_{\tau^{\beta}}\bigg{]}\\ +H^{\tau\alpha}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\mu}H_{\nu\tau}+\bar{\nabla}_{\alpha}\bar{\nabla}_{\nu}H_{\mu\tau}-\bar{\nabla}_{\alpha}\bar{\nabla}_{\tau}H_{\mu\nu}-\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}H_{\tau\alpha}\right)\\ +H_{\mu\nu}\left(\bar{\nabla}_{\alpha}\bar{\nabla}_{\beta}H^{\alpha\beta}-\bar{\nabla}^{2}H\right)+\bar{\nabla}_{\alpha}H^{\alpha\tau}\left(\bar{\nabla}_{\mu}H_{\nu\tau}+\bar{\nabla}_{\nu}H_{\mu\tau}-\bar{\nabla}_{\tau}H_{\mu\nu}\right)\\ -\frac{1}{2}\bar{\nabla}_{\mu}H_{\tau\alpha}\bar{\nabla}_{\nu}H^{\tau\alpha}+\bar{\nabla}_{\alpha}H_{\nu\tau}\bar{\nabla}^{\tau}H_{\mu}^{\alpha}-\bar{\nabla}_{\alpha}H_{\nu\tau}\bar{\nabla}^{\alpha}H_{\mu}^{\tau}\\ -\frac{1}{2}\bar{\nabla}^{\alpha}H\left(\bar{\nabla}_{\mu}H_{\nu\alpha}+\bar{\nabla}_{\nu}H_{\mu\alpha}-\bar{\nabla}_{\alpha}H_{\mu\nu}\right).$ (3)
@@ 3 Solution space at the second order
$\partial_{u}n_{A}=\frac{1}{4}D^{B}D_{A}D^{E}c_{EB}-\frac{1}{4}D^{2}D^{B}c_{BA}-uD_{A}(\partial_{u}m_{b})\\ +\frac{1}{4}D_{A}(C_{BE}N^{BE})-\frac{1}{4}D_{B}(C^{BE}N_{EA})+\frac{1}{2}C_{AB}D_{E}N^{EB},$ (18)
@@ 5 First order supertranslations
$r^{\prime}=r-\frac{G}{2}D^{2}T+\frac{G^{2}}{4r}\bigg{(}D^{A}D^{B}TD_{A}D_{B}T+2D_{A}C^{AB}D_{B}T+C^{AB}D_{A}D_{B}T$
$\quad\quad\quad+2D_{A}TD^{A}T-\frac{1}{2}D^{2}TD^{2}T\bigg{)}+\frac{G^{2}}{6r^{2}}D^{AB}D_{A}D_{B}T$ (33)
@@
when deriving (37). Clearly, the condition of preserving $n_{A}$ in the form of (26) in the new coordinates system will reduce the supertranslation to one translation along time direction $u$, considering that $M_{b}=-\sum{\frac{m}{(v_{\mu}n^{\mu})^{3}}}$ can not be a constant for arbitrary numbers of pointlike bodies. Even for the case of a single particle, $M_{b}$ is not necessary to be a constant and a supertranslation can change the angular momentum of the system which agrees with that the boost charges of the supertranslated Schwarzschild black hole are generically non-vanishing [52]. Note that the velocity of the particle essentially is a boost of the Schwarzschild solution [53].
@@ 6 Concluding remarks
In this paper, we apply the charge conservation to prove that the first order supertranslation should be ruled out in the PM expansion which can serve as the foundation of uniquely determining the angular momentum loss in gravitational scattering. It is worthwhile to point out that ruling out the first order supertranslation only means that the first order supertranslation parameter is fixed, but not necessary to be zero. The precise formula of the first order supertranslation parameter can be obtained for a precise gravitational scattering, such as gravitational Bremsstrahlung [54, 41, 10]. It is amusing to notice that once the first order supertranslation is fixed, the complete first order metric is fixed. The derivation of the other parts of the second order metric are relevant to the first order metric, but can not affect the first order metric. Then the key point of our prescription is that we use part of the higher order solutions in $G$ expansion to fix the lower order supertranslation. It is interesting to see if such idea can be valid at higher order in $G$ to fix the complete supertranslation freedom.
@@
The first piece on the right hand side is the independent second order gauge transformation. The second piece is induced from the first order gauge transformation. It is not a pure gauge transformation at the second order because it will change the first order metric. One can effectively consider that the second order perturbative theory is living on a curved background with metric $\eta_{\mu\nu}+GH_{\mu\nu}$. The effective stress tensor coupled to the second order theory is from curved effect of the background spacetime. Then the first order gauge transformation does change the spacetime metric, hence change the frame of the observer. Note that the translations along spatial dimensions correspond to the shift of the origin of the spacetime which has also been ruled out as they will change the angular momentum. Thus, a first order supertranslation (except the translation along $u$) will set the second order theory to a different frame or a different origin choice than the first order theory. Such interpretation is in consistent with the full Einstein theory [36] where memory effect will change the frame of observer which leads to the supertranslation ambiguity of angular momentum. In the PM expansion, there is no dissipation from the self-part at the second order. The charge conservation can serve as calibration of frames and origin choices of the two observers for describing the first order and second order perturbative theories. The fact that theories at different orders can be in different frames or with different origin choices reflects the non-linear effect from interactions between the gravitational fields and the observers in the full Einstein theory.
@@ Acknowledgments
The authors thank Zhengwen Liu and Kai-Yu Zhang for useful discussions. The authors are grateful to Prof. Massimo Porrati for clarifying several crucial remarks. This work is supported in part by the National Natural Science Foundation of China (NSFC) under Grants No. 11935009 and No. 11905156.
@@ Appendix B Conservation of the effective stress tensor
The conservation of the stress tensor is induced from the Bianchi identity of the Einstein tensor $\nabla_{\mu}G^{\mu}_{\nu}=0$. We consider the Minkowski background and the pointlike bodies source for the first order perturbative theory. Hence the zeroth order of the Einstein tensor vanishes and the first order vanishes outside the trajectory of the pointlike bodies. At the second order, the Bianchi identity outside the trajectory of the pointlike bodies leads to
@@
It is obvious that ¯$\bar{\nabla}^{\mu}E_{\mu\nu}=0$ which can be easily deduced from the fact that the linearized equation $E_{\mu\nu}$ at the second order is invariant under the gauge transformation